# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 5.3 Wall-crossing
where ${\Phi}^{\prime}:{\mathbb{P}}^{1}\to\widehat{C}\overset{{\Phi}}{\to}{\operatorname{Tor}}(P_{\Delta})$ is ${\Phi}$ combined with the normalization, $p^{\prime},p^{\prime\prime}$ are the preimages of $p\in\widehat{C}$. The vanishing follows from the fact that the degree of the line bundle is $-1$.
(6) Suppose that $\xi$ is as in Lemma 5.2(2vi) and that the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{|t-t^{*}|<\zeta}$ has the point $w_{0}(t)$ and two points $w_{1}(t),w_{2}(t)\in\widehat{\boldsymbol{w}}_{\partial}(t)$ in a fixed position, while the other two points $w_{3}(t),w_{4}(t)\in\widehat{\boldsymbol{w}}_{\partial}(t)$ are mobile. Also, for the dimension reason, we can assume that $w_{0}=w_{0}(t^{*})$ is in general position with respect to $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ on the rational curve $C$, which, in particular, yields the following: the point $p_{0}\in\widehat{C}$ is in general position with respect to the remaining points $p_{j}={\Phi}^{-1}(w_{j})\in\widehat{\boldsymbol{p}}_{\partial}\subset\widehat{C}$, $j=1,2,3,4$. By Lemma 2.4, the curve $C$ is immersed and smooth at $\widehat{\boldsymbol{w}}(t^{*})$. We can assume, in addition, that $C$ is nodal (the treatment of more complicated immersed singularities is very similar with a bit more involved notations).
Let $\xi(t)=[{\Phi}_{t}:(\widehat{C}_{t},{\boldsymbol{p}}_{t})\to{\operatorname{Tor}}(P_{\Delta})]\in\overline{\mathcal{M}}_{1,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t))$, $|t-t^{*}|<\zeta$, run over the germ of a real one-dimensional subfamily of {¯$\{\overline{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))\}_{|t-t^{*}|<\zeta}$. It induces a one-dimensional family ${\mathcal{V}}\subset|{\mathcal{L}}_{P_{\Delta}}|$ of curves $C^{(t)}={\Phi}_{t,*}(\widehat{C}_{t})$ on ${\operatorname{Tor}}(P_{\Delta})$, where $C^{(t^{*})}={\Phi}_{*}(\widehat{C})=2C$. Any curve $C^{(t)}$, $t\neq t^{*}$, has four nodes in a neighborhood of each of the $(C^{2}-c_{1}({\operatorname{Tor}}(P_{\Delta})[C])/2+1$ nodes of $C$. We show that the remaining $2c_{1}({\operatorname{Tor}}(P_{\Delta})[C]-4$ nodes of $C^{(t)}$ are located in a neighborhood of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. The intersection of $C^{(t)}$ with a neighborhood of a point $w_{i}(t^{*})$, $1\leq i\leq 4$, is an immersed disc. Following the lines of part (2) of the proof of Lemma 4.3, introduce local conjugation-invariant coordinates $x,y$ in a neighborhood of $w_{j}(t^{*})$ such that $w_{i}(t)=(0,0)$ (which means that, if $w_{i}(t)$ is mobile, then the coordinate system moves too), the toric divisor containing $w_{i}(t)$ is ${\operatorname{Tor}}(\sigma)=\{y=0\}$, and ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})=\{y\geq 0\}$. Then, the equation $F_{t_{1}}(x)=0$ of $C_{t}$, where $t_{1}=t-t^{*}$, is given by formula (14), where we set $k=k_{i}$, and correspondingly, the geometry of $C^{(t)}$ in a neighborhood of $w_{i}(t^{*})$ is described by the real rational affine limit curve $C_{1}$ given by (15). Observe that $C_{1}$ has $k_{i}-1$ nodes and a global real one-dimensional branch having one vertical tangent. We then conclude that
  • the nodes of $C^{(t)}$ are located in a neighborhood of ${\operatorname{Sing}}\hskip 1.70709pt(C)\cup\widehat{\boldsymbol{w}}_{\partial}(t^{*})$;
  • the parameter $\gamma$ in (14) must be even, since otherwise, the change of sign of $t_{1}$ would lead to the change of the position of that global real branch with respect to the vertical tangent, but this position is determined by the double covering ${\Phi}:\widehat{C}\to C$; in particular, it follows that the change of sign of $t_{1}$ does not modify the contribution of the part of $C^{(t)}$ in a neighborhood of $w_{i}(t^{*})$ to formula (9) for $W_{1}(\xi_{t})$.
Let ${\operatorname{Span}}\{2C,C^{\prime}\}$ and ${\operatorname{Span}}\{C^{(t)},C^{\prime}_{t}\}$ be the lines in $|{\mathcal{L}}_{P_{\Delta}}|$ tangent to ${\mathcal{V}}$ at $2C$ and $C^{(t)}$, respectively. For $t\neq t^{*}$, the curve $C^{\prime}_{t}$ intersects $C^{(t)}$ at the point $w_{0}$, at each node, at each point $w_{j}^{\sigma}(t)$ with multiplicity $2k_{j}^{\sigma}$ or $12k_{j}^{\sigma}-$ 1 according as $w_{j}^{\sigma}$ is fixed or mobile, and at one more point $w^{\prime}$ (follows from the Bézout theorem). Lifting these points to $\widehat{C}^{(t)}$ and passing to the limit as $t\to t^{*}$, we obtain on $\widehat{C}$ the divisor
$D=p_{0}+4{\Phi}^{*}({\operatorname{Sing}}\hskip 1.70709pt(C))+(4k_{1}-2)p_{1}+(4k_{2}-2)p_{2}+(4k_{3}-3)p_{3}+(4k_{4}-3)p_{4}+p^{\prime}\ .$
Due to the general position of $w_{0}$ on $C$, we get $p^{\prime}\not\in\widehat{\boldsymbol{p}}_{\partial}\cup\{p_{0}\}$. We use this to show that for $t^{\prime}<t^{*}<t^{\prime\prime}$, the points $w_{i}(t^{\prime})$ and $w_{i}(t^{\prime\prime})$ are separated on the toric divisor by the point $w_{i}(t^{*})$, $i=3,4$. Indeed, otherwise, we would get $w_{i}(t^{\prime})=w_{i}(t^{\prime\prime})$ for, say, $i=3$ and for $t^{\prime}<t^{*}<t^{\prime\prime}$ depending on each other and converging to $t^{*}$. Lifting the intersection $C^{(t^{\prime})}\cap C^{(t^{\prime\prime})}$ upon $\widehat{C}_{t^{\prime}}$, we would obtain on $\widehat{C}$ the divisor
$D^{\prime}=p_{0}+4{\Phi}^{*}({\operatorname{Sing}}\hskip 1.70709pt(C))+(4k_{1}-2)p_{1}+(4k_{2}-2)p_{2}+(4k_{3}-2)p_{3}+(4k_{4}-3)p_{4}\ ,$
not linearly equivalent to $D$, which is a contradiction. Thus, the real part of the family ${\mathcal{V}}$ homeomorphically projects onto the interval $|t-t^{*}|<\zeta$, and the sign $W_{1}(\xi(t))$ remains constant as $0<|t-t^{*}|<\zeta$. $\Box$
@@
We closely follow the argument in the proof of Theorems 3.2 and 3.7. Take two sequences ${\boldsymbol{w}}(0)=({\boldsymbol{w}}_{\partial}(0),w_{1}(0),w_{2}(0))$ and ${\boldsymbol{w}}(1)=({\boldsymbol{w}}_{\partial}(1),w_{1}(1),w_{2}(t))$, satisfying the conditions of Theorem 3.7. We may assume that $\widehat{\boldsymbol{w}}_{\partial}(0),\widehat{\boldsymbol{w}}_{\partial}(1)\in\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ for some $\tau>0$. Then, we join these sequences by a generic path $\{\widehat{\boldsymbol{w}}_{\partial}(t)\}_{0\leq t\leq 1}$, in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ and the points $w_{i}(0)$, $w_{i}(1)$ by a generic path $\{w_{i}(t)\}_{0\leq t\leq 1}$ in the respective open quadrant, $i=1,2$. Then we verify the constancy of $W_{2}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ inside the chambers (i.e., connected components of $\operatorname{Men}_{2}(\Delta)$) and in all possible wall-crossing events, for all $\kappa$.
@@ 6.1 Invariance inside chambers
Let us verify the constancy of $W_{2}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ along intervals $t^{\prime}<t<t^{\prime\prime}$ such that ${\boldsymbol{w}}(t)\in\operatorname{Men}_{2}(\Delta)$ for all $t^{\prime}<t<t^{\prime\prime}$. To this end, it is enough to show that the projection ${\mathcal{M}}_{2,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}\to(t^{\prime},t^{\prime\prime})$ has no critical points (cf. the proof of Theorems 3.2 and 3.7). This requirement is a consequence of the following statement similar to Lemmas 4.1 and 5.1 ( cf. [23, Lemma 2.3]).
@@
Let $\xi=[{\Phi}:(S_{2},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{2,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}})$, where $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{2}(\Delta)$. Then
$H^{0}\left(S_{2},{\mathcal{N}}_{\Phi}\left(-p_{1}-p_{2}-\sum_{s=0}^{2}\sum_{i=1}^{n_{s}}2k_{i}^{(s)}p_{i}^{(s)}+p_{n_{0}}^{(0)}\right)\right)=0,$ (27)
where $[{\Phi}:(S_{2},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in{\mathcal{M}}_{2,n}^{\mathbb{R}}({\mathbb{P}}^{2},\Delta,\widehat{\boldsymbol{w}}(t))$.
Proof. The non-vanishing in (27) would mean the existence of a real plane curve $C^{\prime}\neq C={\Phi}(S_{2})$ of degree $2d$ passing through $w_{1}$ and $w_{2}$, intersecting $D_{s}$, $s=0,1,2$, at each point $n_{i}^{(s)}$, $1\leq i\leq n_{s}$, $(s,i)\neq(0,n_{0})$, with multiplicity $2k_{i}^{(s)}$, intersecting $D_{0}$ at $w_{n_{0}}^{(0)}$ with multiplicity $12k_{n_{0}}^{(0)}-$ 1, and intersecting $C$ at each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)$ with multiplicity $\geq\delta(C,z)$. Note that $C^{\prime}$ must additionally intersect $C$ in an extra point on the real branch in $({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}$ and in an extra point on the real branch in ()±$({\mathbb{P}}_{\mathbb{R}})^{2}_{\pm}$. However, this leads to a contradiction with the Bézout theorem:

### v2

@@ 5.3 Wall-crossing
where ${\Phi}^{\prime}:{\mathbb{P}}^{1}\to\widehat{C}\overset{{\Phi}}{\to}{\operatorname{Tor}}(P_{\Delta})$ is ${\Phi}$ combined with the normalization, $p^{\prime},p^{\prime\prime}$ are the preimages of $p\in\widehat{C}$. This $h^{1}$-vanishing follows from the fact that the degree of the line bundle is
$(-CK_{{\operatorname{Tor}}(P_{\Delta})}-2)+(CK_{{\operatorname{Tor}}(P_{\Delta})}-2k_{i}^{\sigma}+2)-((2l_{1}+1)+(2l_{2}+1)+1)=-1.$
(6) Suppose that $\xi$ is as in Lemma 5.2(2vi). Since the rational curve $C={\Phi}(\widehat{C})$ has one global one-dimensional real branch and contains the point $w_{0}\in Q$, it intersects the toric divisor ${\operatorname{Tor}}(\sigma)$ at two points $w_{1},w_{2}$ with odd multiplicities $k_{1},k_{2}$, respectively, where ${\operatorname{Tor}}^{+}_{\mathbb{R}}(\sigma)$ is the (unique) common side of ${\operatorname{Tor}}^{+}_{\mathbb{R}}(P_{\Delta})$ and $\overline{Q}$. Correspondingly, $P_{\Delta}$ is a lattice triangle, and each of the two other toric divisors intersects $C$ with an even multiplicity, see Figure 7(a) (the two fragments of ${\mathbb{R}}C$ shown in bold are doubly covered by the two ovals of ${\mathbb{R}}\widehat{C}$).
[…]
Figure 7: Proof of Lemma 4.3, part (6)
We define the path $\{(\widehat{\boldsymbol{w}}_{\partial}(t),w_{0}(t))\}_{|t-t^{*}|<{\varepsilon}}$ by fixing the position of $w_{1}$, $w_{2}$, $w_{3}$, and $w_{4}$ and leaving $w_{0}$ mobile. Note that $w_{0}$ can be chosen in general position on $C$. We suppose that $\{w_{0}(t)\}_{|t-t^{*}|<{\varepsilon}}$ sweeps the germ of a real line $L\subset Q$ transversally intersecting ${\mathbb{R}}C$ at $w_{0}=w_{0}(t^{*})$.
Note also that $(C,(w_{0},w_{1},w_{2},w_{3},w_{4}))$ has two preimages: two marked curves $(\widehat{C},(p^{\prime}_{0},p_{1},p_{2},p_{3},p_{4}))$ and $(\widehat{C},(p^{\prime\prime}_{0},p_{1},p_{2},p_{3},p_{4}))$. It follows from (8) that the quantum index of the two latter curves equipped with any complex orientation orientation is zero. The signs (10) and (11) are preserved too when we reverse the orientation.
The family {¯$\{\overline{\mathcal{M}}_{1,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))\}_{|t-t^{*}|<{\varepsilon}}$ naturally projects to a bunch of real one-dimensional subvarieties in the linear system $|{\mathcal{L}}_{P_{\Delta}}|$. Denote by ${\mathcal{V}}$ the union of those components of the image that pass through $2C$ and have the same tangent line (${\mathcal{V}}$ is determined up to a finite choice). Let this common tangent line $\Lambda$ be spanned by $2C$ and some $C^{*}\in|{\mathcal{L}}_{P_{\Delta}}|\setminus\{2C\}$. Two cases are possible: either $C^{*}$ contains $C$ as a component, or not.
Let $C\not\subset C^{*}$. Introduce local conjugation-invariant coordinates $x,y$ in a neighborhood of $w_{0}$ so that
$w_{0}=(0,0),\ L=\{y=0\},\ C=\{x=0\},\ C^{*}=\left\{a+\sum_{i+j>0}a_{ij}x^{i}y^{j}=0\right\},$
where $a\neq 0$. Let us parameterize $\Lambda$ as follows:
$C(\theta)=\left\{x^{2}+\theta\left(a+\sum_{i+j>0}a_{ij}x^{i}y^{j}\right)=0\right\},\quad\theta\in({\mathbb{R}},0).$
Then, for $a\theta<0$, the curve $C(\theta)$ intersects $L$ at two points, $x=\sqrt{-a\theta}+o(\sqrt{\theta})$ and $x=-\sqrt{-a\theta}+o(\sqrt{\theta})$. The same holds for any curve $C^{\prime}\in{\mathcal{V}}$ projecting to $C(\theta)$; hence, $C^{\prime}$ contributes to the count of $W_{1}^{0}(t)$ and $\widetilde{W}_{1}^{0}(t)$ both, for $t<t^{*}$ and for $t>t^{*}$. This yields the constancy of $W_{1}^{0}(t)$ and $\widetilde{W}_{1}^{0}(t)$ over the segment $t\in(t^{*}-{\varepsilon},t^{*}+{\varepsilon})$.
Let $C^{*}=C\cup C^{\prime}$, $C^{\prime}\neq C$. In the above coordinates $x,y$, we have
$C^{*}=\left\{x\left(a+\sum_{i+j>0}a_{ij}x^{i}y^{j}\right)=0\right\},\quad a\neq 0.$
Thus,
$\Lambda=\left\{C(\theta):=\left\{x\left(x+\theta\left(a+\sum_{i+j>0}a_{ij}x^{i}y^{j}\right)\right)=0\right\}\right\}_{\theta\in{\mathbb{R}},0)},$
$C(\theta)\cap L=\{w_{0},w_{0}(t)\},\quad w_{0}(t)=(-a\theta+O(\theta^{2}),0).$ (41)
Without loss of generality, we can assume that ${\mathcal{V}}$ consists of one component. If ${\mathcal{V}}$ homeomorphically projects onto the germ $(\Lambda,2C)$, and hence onto the segment $(t^{*}-{\varepsilon},t^{*}+{\varepsilon})$, we establish the constancy of $W_{1}^{0}(t)$ and $\widetilde{W}_{1}^{0}(t)$ by showing that a curve $C^{(t)}$ makes the same contribution to $W_{1}^{0}(t)$ both for $t<t^{*}$ and for $t>t^{*}$, and so for $\widetilde{W}_{1}^{0}(t)$.
To this end, we analyze the singularities of curve $C^{(t)}$, $0<|t-t^{*}|<{\varepsilon}$. The curve $C$ has $(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2)/2$ nodes. In a neighborhood of each of these nodes, the curve $C^{9t)}$ has four nodes: four hyperbolic nodes in a neighborhood of a hyperbolic node of $C$, two elliptic and two complex conjugate nodes in a neighborhood of an elliptic node of $C$, four non-real nodes in a neighborhood of a non-real node of $C$. There remain
$\frac{4C^{2}+2CK_{{\operatorname{Tor}}(P_{\Delta})}}{2}-4\cdot\frac{C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2}{2}=k_{1}+k_{2}+k_{3}+k_{4}-4$
nodes of $C^{(t)}$. We show that they are located in a neighborhood of the points $w_{1},w_{2},w_{3},w_{4}$. Consider the tropical limit of the family of the parameterized curves $\{{\Phi}_{t}:\widehat{C}^{(t)}\to C^{(t)}\hookrightarrow{\mathfrak{X}}_{t}={\operatorname{Tor}}(P_{\Delta})\}_{C^{(t)}\in{\mathcal{V}},0<|t-t^{*}|<{\varepsilon}}$ as $t\to t^{*}$ (see details in [6, Section 1], where the setting is most suitable for our argument). The limit of the maps ${\Phi}_{t}:\widehat{C}^{(t)}\to{\mathfrak{X}}_{t}={\operatorname{Tor}}(P_{\Delta})$, $t\to t^{*}$, is ${\Phi}_{0}:\widehat{C}^{(0)}\to{\mathfrak{X}}_{0}$, where $\widehat{C}^{(0)}$ is the union of the elliptic curve $\widehat{C}$ with four marked points $p_{1},p_{2},p_{3},p_{4}$ mapped to $w_{1},w_{2},w_{3},w_{4}$, respectively, and of connected curves $\widehat{C}_{1},\widehat{C}_{2},\widehat{C}_{3},\widehat{C}_{4}$ of arithmetic genus zero attached to $p_{1},p_{2},p_{3},p_{4}$, respectively. Pick $m\in\{1,2,3,4\}$ and introduce local conjugation-invariant coordinates $x_{m},y_{m}$ in a neighborhood of $w_{m}$ so that $w_{m}=(0,0)$, the toric divisor passing through $w_{m}$ is given by $\{y=0\}$, and the germ of $C$ at $w_{m}$ is given by $\{y-x^{k_{m}}=0\}$. In particular, a local equation of $2C$ reads
$(y_{m}-x^{k_{m}})^{2}=0.$
Components of $\widehat{C}_{m}$ are mapped to toric components of ${\mathfrak{X}}_{0}$ that correspond to a certain subdivision of the region under the segment $[(0,2),(2k_{m},0)]$. In view of the given data, the only possible subdivision contains just one triangle $T_{m}={\operatorname{conv}}\{(0,1),(0,2),(2k_{m},0)\}$ (see Figure 7(b)). The limit curve $C_{m}\subset{\operatorname{Tor}}(T_{m})$ is rational, quadratically tangent to the toric divisor ${\operatorname{Tor}}([(0,2),(2k_{m},0)])$, and it has $k_{m}-1$ nodes (cf. [28, Lemma 3.5]). An equation of $C_{m}$ reads
$(y_{m}-x_{m}^{k_{m}})^{2}+\sum_{i=0}^{k_{m}-1}a_{i}^{(m)}x_{m}^{i}y_{m}=0,\quad a_{0}^{(m)}\neq 0.$
Correspondingly, the curves $C^{(t)}\in{\mathcal{V}}$ are given in a neighborhood of $w_{m}$ by
[…]
where $\tau=t-t^{*}$, and h.o.t. denotes terms with higher powers of $\tau$ in the preceding monomials as well as monomials above the segment $[(0,2),(2k_{m},0)]$.
If $m=1$ or $2$, then $k_{m}$ is odd. In this case, all exponents of $\tau$ in the second summand of (42) must be even since otherwise the coefficient of $y_{m}$ will change sign when $\tau$ does, but geometrically this would mean that the local real branch of $C^{(t)}$ at $w_{m}$ would jump from the quadrant $Q$ to the positive quadrant when $t$ transits through $t^{*}$ which would contradict the fact that $C^{(t)}$ must have a global real branch in $Q$. This observation yields that the distribution among the quadrants of the elliptic nodes in a neighborhood of $w_{m}$ does not change when $\tau$ changes its sign. If $m=3$ or $4$, then $k_{m}$ is even. If all exponents of $\tau$ in the second summand of (42) are even, we obtain the preceding conclusion. If there are odd exponens, then the change of sign of $\tau$ geometrically is equivalent to the reflection of the limit curve $C_{m}$ with respect to the $y_{m}$-axis, which again does not affect the distribution of elliptic nodes among quadrants. In all the cases, we immediately derive the constancy of $W_{1}^{0}(t)$ and $\widetilde{W}_{1}^{0}(t)$ along the segment $t\in(t^{*}-{\varepsilon},t^{*}+{\varepsilon})$.
It remains to show that the case of ${\mathcal{V}}$ doubly covering a half of the interval $(t^{*}-{\varepsilon},t^{*}+{\varepsilon})$ with ramification at $t^{*}$ is not possible. We argue by contradiction based on the Bézout theorem. Let $C^{(t^{\prime})},C^{(t^{\prime\prime})}\in{\mathcal{V}}$ be two curves with $t^{\prime}<t^{*}<t^{\prime\prime}$ that intersect the line $L$ at the same point $w_{0}(t^{\prime})=w_{0}(t^{\prime\prime})$ given by the second formula in (41). In a neighborhood of ${\operatorname{Sing}}\hskip 1.70709pt(C)$, the curves $C^{(t^{\prime})}$ and $C^{(t^{\prime\prime})}$ have at least
$8\cdot\frac{C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}+2}{2}=4C^{2}+4CK_{{\operatorname{Tor}}(P_{\Delta})}+8=4C^{2}-\sum_{m=1}^{4}(4k_{m}-2)$
intersection points. To estimate the intersection multiplicity of $C^{(t^{\prime})}$ and $C^{(t^{\prime\prime})}$ in a neighborhood of $w_{m}$, $m=1,2,3,4$, we rescale the formula (42) bringing it to the common for $t^{\prime}$ and $t^{\prime\prime}$ form (since $|t^{\prime}-t^{*}|$ and $|t^{\prime\prime}-t^{*}|$ are asymptotically equivalent as $t^{\prime},t^{\prime\prime}\to t^{*}$)
$(y_{m}-x_{m}^{k_{m}})^{2}+\sum_{i=0}^{k_{m}-1}({\operatorname{sign}}\tau)^{(k_{m}-i)r_{m}}a_{i}^{(m)}x_{m}^{i}y_{m}+O(\tau)=0.$
In a neighborhood of $w_{m}$, $m=1,2,3,4$, the curves $C^{(t^{\prime})},C^{(t^{\prime\prime})}$ intersect with multiplicity at least
$24k_{m}-$ 2 .
Indeed, if the change of sign of $\tau$ in (42) yields the same limit curve $C_{m}$, then $C^{(t^{\prime})}$ and $C^{(t^{\prime\prime})}$ intersect with multiplicity at least $2k_{m}$ at $w_{m}$ and intersect with multiplicity at least $2(k_{m}-1)$ near the nodes of $C_{m}$. If the change of sign of $\tau$ induces the reflection of $C_{m}$ against the $y_{m}$-axis, then $C^{(t^{\prime})}$ and $C^{(t^{\prime\prime})}$ intersect in the big torus of ${\operatorname{Tor}}(T_{m})$ with multiplicity $22k_{m}-$ 2 (here we subtract the intersection multiplicity at the common tangency point on the toric divisor ${\operatorname{Tor}}([(0,2),(2k_{m},0])$) and they intersect at $w_{m}$ with multiplicity $\geq 2k_{m}$. Taking into account the extra intersection point $w_{0}(t^{\prime})=w_{0}(t^{\prime\prime})\in L$, we arrive to a contradiction with the Bézout theorem. $\Box$
@@
We closely follow the argument in the proof of Theorems 3.2 and 3.6. Take two sequences ${\boldsymbol{w}}(0)=({\boldsymbol{w}}_{\partial}(0),w_{1}(0),w_{2}(0))$ and ${\boldsymbol{w}}(1)=({\boldsymbol{w}}_{\partial}(1),w_{1}(1),w_{2}(t))$, satisfying the conditions of Theorem 3.6. We may assume that $\widehat{\boldsymbol{w}}_{\partial}(0),\widehat{\boldsymbol{w}}_{\partial}(1)\in\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ for some $\rho>0$. Then, we join these sequences by a generic path $\{\widehat{\boldsymbol{w}}_{\partial}(t)\}_{0\leq t\leq 1}$, in $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ and the points $w_{i}(0)$, $w_{i}(1)$ by a generic path $\{w_{i}(t)\}_{0\leq t\leq 1}$ in the respective open quadrant, $i=1,2$. Then we verify the constancy of $W_{2}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ inside the chambers (i.e., connected components of $\operatorname{Men}_{2}(\Delta)$) and in all possible wall-crossing events, for all $\kappa$.
@@ 6.1 Invariance inside chambers
Let us verify the constancy of $W_{2}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ along intervals $t^{\prime}<t<t^{\prime\prime}$ such that ${\boldsymbol{w}}(t)\in\operatorname{Men}_{2}(\Delta)$ for all $t^{\prime}<t<t^{\prime\prime}$. To this end, it is enough to show that the projection ${\mathcal{M}}_{2,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}\to(t^{\prime},t^{\prime\prime})$ has no critical points (cf. the proof of Theorems 3.2 and 3.6). This requirement is a consequence of the following statement similar to Lemmas 4.1 and 5.1 (cf. [29, Lemma 2.3]).
@@
Let $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in{\mathcal{M}}_{2,n}^{\;{\mathbb{R}}}(\Delta,\widehat{\boldsymbol{w}})$, where $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{2}(\Delta)$. Then,
$h^{0}\left(\widehat{C},{\mathcal{N}}_{\Phi}\left(-p_{1}-p_{2}-\sum_{s=0}^{2}\sum_{i=1}^{n_{s}}2k_{i}^{(s)}p_{i}^{(s)}+p_{n_{0}}^{(0)}\right)\right)=0,$ (43)
where $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in{\mathcal{M}}_{2,n}^{\mathbb{R}}({\mathbb{P}}^{2},\Delta,\widehat{\boldsymbol{w}}(t))$.
Proof. The non-vanishing in (43) would mean the existence of a real plane curve $C^{\prime}\neq C={\Phi}(\widehat{C})$ of degree $2d$ passing through $w_{1}$ and $w_{2}$, intersecting $D_{s}$, $s=0,1,2$, at each point $n_{i}^{(s)}$, $1\leq i\leq n_{s}$, $(s,i)\neq(0,n_{0})$, with multiplicity $2k_{i}^{(s)}$, intersecting $D_{0}$ at $w_{n_{0}}^{(0)}$ with multiplicity $12k_{n_{0}}^{(0)}-$ 1, and intersecting $C$ at each point $z\in{\operatorname{Sing}}\hskip 1.70709pt(C)$ with multiplicity $\geq\delta(C,z)$. Note that $C^{\prime}$ must additionally intersect $C$ in an extra point on the real branch in $({\mathbb{P}}_{\mathbb{R}}^{2})^{(-,+)}$ and in an extra point on the real branch in $({\mathbb{P}}_{\mathbb{R}}^{2})^{(+,-)}$. However, this leads to a contradiction with the Bézout theorem: